According to the Environmental Defense Fund, “Americans are eating more and more salmon, drawn to its rich taste and health benefits. Increasingly they are choosing farmed salmon because of its wide availability and low price. But in the last few years, farmed salmon has been surrounded by controversy over its health risks and the ecological impacts of salmon aquaculture operations. Studies have shown that some farmed salmon is relatively higher in contaminants like PCBs than wild salmon, and there is mounting concern over the industry’s impact on wild salmon populations.” In a widely cited study of contaminants in farmed salmon, fish from many sources were analyzed for 14 organic contaminants. One of those was the insecticide mirex, which has been shown to be carcinogenic and is suspected of being toxic to the liver, kidneys, and endocrine system. Summaries for 150 mirex concentrations (in parts per million) from a variety of farmed salmon sources were reported as: n = 150; x ̅ = 0.0913ppm; s = 0.0495ppm The Environmental Protection Agency (EPA) recommends to recreational fishers as a “screening value” that mirex concentrations be no larger than 0.08 ppm. What does the 95% confidence interval say about that value?
Answer #1

Given : = 0.0913, s = 0.0495, n = 150, = 0.05

Although population standard deviation is unknown but n < 30, the tcritical (2 tail) for = 0.05, , is 1.96

The Confidence Interval is given by ME, where

The Lower Limit = 0.0913 - 0.0067 = 0.0846

The Upper Limit = 0.0913 + 0.0067 = 0.0980

The 95% Confidence Interval is 0.0846 < < 0.0980

From this interval we see that the value 0.0846 is the lower limit, which means that mirex concentrations in the farmed salmon are greater than what is recommended by the EPA to recreational fishermen.